OpenStudy (anonymous):

f(x) = x - 1/6X^2 - 2/3lnX a) List the Vertical and horizontal asymptote. b) The intervals of increase and decrease. c) The local max and min values. d) Intervals of concavity. Please help!

OpenStudy (anonymous):

is the function this? \[\huge f(x)=x-\frac{1}{6}x^2-\frac{2}{3}lnx \] or is it \[\huge f(x)=x-\frac{1}{6x^2}-\frac{2}{3lnx} \]

OpenStudy (dumbcow):

http://www.wolframalpha.com/input/?i=x-%281%2F6%29x^2-%282%2F3%29log%28x%29 asymptote at x=0 for increasing/decreasing look at when derivative is either pos/neg for concavity look at when 2nd derivative is pos/neg